OpenStudy (anonymous):

Solve 0≤θ<π 2sin2θ=3(cos2θ+sinθ−1)

OpenStudy (experimentx):

Let \( \sin(x) = u \) then \( \cos (x) = \sqrt{1-u^2}\) \[ 2 \sin 2\theta = 3 ( \cos 2\theta + \sin \theta - 1) \\ 4 u \sqrt{1 - u^2} = 3(1 - 2u^2 + u -1)\] solve this equal for \( u \)

OpenStudy (anonymous):

I managed to get x=\[\sin^{-1} \left( \sqrt{\frac{ 1 }{ 6 }} \right)\] Which then gave me x=24 is this correct?

OpenStudy (experimentx):

x=0, is for sure from here http://www.wolframalpha.com/input/?i=Solve+4+x+sqrt%281-x%5E2%29+%3D+3+%28+1+-+2x%5E2+%2B+x-1%29 the other one is beyond the given range

OpenStudy (anonymous):

Thanks, I just hope this doesn't come up in the exam as it was quite time consuming